OpenStudy (anonymous):

Please explain in words the meaning of this notation. (f ° g)(3)

OpenStudy (anonymous):

f of g of x f(g(x)) plug x into g(x) and solve, place that solved value as the input in f(x) and solve for final answer

OpenStudy (anonymous):

x in this case is 3 so solve f(3) 1st

OpenStudy (anonymous):

@ehuman is there anything else ??

OpenStudy (anonymous):

(f ° g)(3) means you take the value of 3 in place of x in G(x) then use that value in place of x and solve f(x)

OpenStudy (jdoe0001):

(f of g(x)) or as the notation goes (fog)(x) is just g(x) as argument of the function f(x) so f(x) is a HOST function to the CLIENT g(x) let's say for example, let us use these 2 functions \(\bf f(x) = x^2+2x-1\qquad \qquad g(x) = \color{red}{5 - x}\\ \quad \\ (fog)(x)= f(\quad g(x)\quad )\implies (\quad \color{red}{g(x)}\quad )^2+2(\quad \color{red}{g(x)}\quad )-1\\ \quad \\ (fog)(3)= f(\quad g(3)\quad )\implies (\quad \color{red}{g(3)}\quad )^2+2(\quad \color{red}{g(3)}\quad )-1\\ \quad \\ \color{blue}{g(3) = 5 - (3)\implies 2}\qquad thus\\ \quad \\ (fog)(3)= f(\quad g(3)\quad )\implies (\quad \color{red}{g(3)}\quad )^2+2(\quad \color{red}{g(3)}\quad )-1\\ \quad \\ \implies (fog)(3)= f(\quad g(3)\quad )\implies (\quad \color{red}{2}\quad )^2+2(\quad \color{red}{2}\quad )-1\)